OpenStudy (anonymous):

Write the equation of a line with the slope of -2 which contains the point (3,6)

OpenStudy (luigi0210):

Use the point slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\]

OpenStudy (luigi0210):

Where m=slope and x1/y1 are (x, y)

OpenStudy (anonymous):

Plug value of the slope and point ( -3 , 6 ) in Y = m x + b 6.00 = 3 + b b= 6 -3 b= 3 So the equation will be Y = -1 x + 3

OpenStudy (texaschic101):

slope is -2, not -1 Luigi

OpenStudy (texaschic101):

y = mx + b (3,6) x = 3 and y = 6 slope(m) = -2 now sub all of that into y = mx + b and solve for b

OpenStudy (luigi0210):

@texaschic101 what method is that?

OpenStudy (texaschic101):

you do not have to use point slope formula if it doesn't ask for point slope formula in the answer...you can just as easily use slope intercept formula

OpenStudy (luigi0210):

Alrighty

OpenStudy (texaschic101):

y = mx + b 6 = -2(3) + b 6 = -6 + b 6 + 6 = b 12 = b equation is : y = -2x + 12 or y - y1 = m(x - x1) y - 6 = -2(x - 3) y - 6 = -2x + 6 y = -2x + 6 + 6 y = -2x + 12 you will get the same answer either way. I just find it easier to use slope intercept, rather then point slope, unless it asks for point slope.